lifewmylilweeb:

If the distance between two objects is decreased to 1 10 of the original distance, how will it change the force of attraction between them? A. The new force will be 100 times more than the original. B. The new force will be 20 times more than the original. C. The new force will be 1/20 of the original. D. The new force will be 1/100 of the original. E. The new force will be 1/10 of the original.

lifewmylilweeb:

please hurry if you can

minustempo:

I believe it is similar to this: Imagine you have a square flashlight. It produces the same amount of light no matter where you shine it. Now imagine you shine it on a square floor, and the light only hits 1 square. how many squares would you cover if you were 3 times as far away? You'd cover 9 squares.

minustempo:

I believe the denominator in the force equation has distance^2 as the denominator. Use that

lifewmylilweeb:

@minustempo wrote:
I believe the denominator in the force equation has distance^2 as the denominator. Use that
im confused so uhm how do i solve the problem

minustempo:

So pretend the numerator is 1. If the distance is 10, then the force is 1/10^2 = 1/100 but if the distance is 1, the force is just 1/1^2 = 1

minustempo:

Decreasing distance to 1/10 increases force 100 fold

lifewmylilweeb:

oooooooooooooooo

minustempo:

yeahhhhhhhhhhhhh

minustempo:

wow that was exactly the same length

lifewmylilweeb:

thank you

minustempo:

you're welcome